MCQ1. The chart below shows the prices on two consecutive days for Wal-Mart, a component stock of the Dow Jones Industrial Average. The chart also shows the sum of the 30 stock prices on each of the days. Use this price information to answer the following questions. (Hint: the PWI is calculated as the adjusted average price) Stock Close: t Close: t+1 Wal-Mart (WMT) 49.6 24.47 30 sum of Prices ( 1505.74 1469.29 i: 0 The Dow Jones Industrial Average closed at a value of 12,240.01 on day t, based on a divisor of 0.1230179. On day t+1, Wal-Mart (WMT) experiences a 2-for-1 stock split. Calculate the return on the Dow Jones Industrial Average from day t to day t+1.
Answer #1

At time T:

The Dow Jones Industrial Average closed at 12,240.01, based on a divisor of 0.1230179

Sum of Prices of 30 stocks = 1505.74

Calculation of Divisor = sum of prices of 30 stocks/ Dow Jones Industrial Average

= 1505.74/ 12,240.01 = 0.12301787 = 0.1230179 (approx.)

At time T+1:

Sum of Prices of 30 stocks = 1469.29

Divisor = sum of prices of 30 stocks/ Dow Jones Industrial Average

0.1230179 = 1469.29/ Dow Jones Industrial Average

Dow Jones Industrial Average = 1469.29/ 0.1230179

= 11,943.709

Calculation of Return on Dow Jones Industrial Average (DJIA) from T to T+1:

Return on DJIA = (DJIA at T+!) - (DJIA at T)/ DJIA at T

= (11,943.709 - 12,240.01)/ 11,943.709

= -296.301/ 11,943.709

= -0.02480812

= -2.48% (approx)